OpenStudy (anonymous):

Medal****Please Help Given the vectorfield F(x,y)= (x+2xsiny) i + (x^2cosy +2y) j Determine whether F is conservative. If it is, find a potential function

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Hint : curl must be 0 for the vector field to be conservative

OpenStudy (anonymous):

what formula would i apply here?

ganeshie8 (ganeshie8):

\(F = Mi + Nj\) \(\text{curl(F)} = N_x - M_y\)

ganeshie8 (ganeshie8):

F(x,y)= (x+2xsiny) i + (x^2cosy +2y) j M = x + 2xsiny N = x^2cosy + 2y N_x = ? M_y = ?

OpenStudy (anonymous):

i take the derivative?

ganeshie8 (ganeshie8):

N_x means the partial of N with respect to x

OpenStudy (anonymous):

2xcosy?

ganeshie8 (ganeshie8):

what is 2xcosy ?

OpenStudy (anonymous):

Nx

ganeshie8 (ganeshie8):

OK, work the curl

OpenStudy (anonymous):

2xcosy-My....i dont know the My

OpenStudy (loser66):

.

ganeshie8 (ganeshie8):

You need to work M_y too

OpenStudy (anonymous):

-2xsiny+2 for My?

ganeshie8 (ganeshie8):

M = x + 2xsiny M_y = ?

OpenStudy (anonymous):

1+2cosy

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

2xcosy

ganeshie8 (ganeshie8):

Yes, curl = ?

OpenStudy (anonymous):

2xcosy-2xcosy=0

ganeshie8 (ganeshie8):

Since the curl is 0, the given vector field is conservative and a potential function exists

ganeshie8 (ganeshie8):

Let \(f(x,y)\) be the potential function, then this must satisfy : \(f_x = M\) \(f_y = N\)

OpenStudy (anonymous):

what if it wasnt 0, it would be non-conservative but still a potential function?

ganeshie8 (ganeshie8):

If the curl is not 0, then the vector field is not conservative and consequently there will not be a potential function. You cannot find a potential function.

ganeshie8 (ganeshie8):

potential function for a vector field exists if and only if the curl of vector field is 0

OpenStudy (anonymous):

oki! from fx=m fy=n....whats the next step?

ganeshie8 (ganeshie8):

plug in m and n

ganeshie8 (ganeshie8):

\[f_x = x + 2x\sin y\] simply integrate both sides with respect to \(x\) to get \(f\)

OpenStudy (anonymous):

x^2/2 +2x^3/3siny?

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

and what about the integration constant ?

OpenStudy (anonymous):

x.... x^2/2+x^2siny

OpenStudy (anonymous):

x^2/2+2x^2/2siny

ganeshie8 (ganeshie8):

Easy, \[f_x = x + 2x\sin y\] integrating both sides with respect to \(x\) gives \[f = \frac{x^2}{2} + x\sin y + \color{red}{g(y)}\]

ganeshie8 (ganeshie8):

that \(\color{red}{g(y)}\) is the arbitrary constant, it shows up everytime you integrate

ganeshie8 (ganeshie8):

you need to find \(\color{red}{g(y)}\)

OpenStudy (anonymous):

xsiny why isnt it x^2siny

ganeshie8 (ganeshie8):

Oops! my mistake... see if below looks fine \[f_x = x + 2x\sin y\] integrating both sides with respect to \(x\) gives \[f = \frac{x^2}{2} + x^2\sin y + \color{red}{g(y)}\]

OpenStudy (anonymous):

i take the third derviative to find g(Y)

ganeshie8 (ganeshie8):

There is an easy way take the derivative of \(f\) with respect to \(y\) and compare it with the equation \(f_y = N\)

ganeshie8 (ganeshie8):

\[f = \frac{x^2}{2} + x^2\sin y + \color{red}{g(y)}\] \[f_y = ?\]

OpenStudy (anonymous):

Y^2

ganeshie8 (ganeshie8):

how

OpenStudy (anonymous):

nevermind...is it x^2cosy

ganeshie8 (ganeshie8):

\[f = \frac{x^2}{2} + x^2\sin y + \color{red}{g(y)}\] \[f_y = x^2\cos y + \color{red}{g'(y)}\]

ganeshie8 (ganeshie8):

compare this with the other equation \(f_y = N\)

OpenStudy (anonymous):

g'(y) is 2y?

ganeshie8 (ganeshie8):

Yes, integrate and solve \(g(y)\)

OpenStudy (anonymous):

y^2

ganeshie8 (ganeshie8):

Looks good! plug that in the potential function and you're done!

ganeshie8 (ganeshie8):

\[f = \frac{x^2}{2} + x^2\sin y + \color{red}{g(y)} = \frac{x^2}{2} + x^2\sin y + \color{red}{y^2} \]

OpenStudy (anonymous):

thank you so much! x^2/2+x^2siny+Y^2

ganeshie8 (ganeshie8):

just a sanity check : find the partials and see if you really get M and N

ganeshie8 (ganeshie8):

you must get \(f_x = M\) and \(f_y = N\)

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

thank you again ganeshie8!!

OpenStudy (irishboy123):

for completeness, there's a really handy way around the final step that avoids fiddling around with the integration constants. because the field is conservative, it is path independent so you can do a line integral. \(W = \int \vec F \bullet \vec dr = \int <x+2xsiny, x^2cosy +2y> \bullet <dx, dy>\) from (0,0) to (x,y) using, for example, these steps:  the line integral is straightforward

ganeshie8 (ganeshie8):

I prefer the line integral too (ofcourse when eyeballing method fails :))